OpenStudy (anonymous):

if 5^x + 4^y = 3189 and 5^(x-3) + 4^(y+1) =281 then x-y= ?

OpenStudy (apoorvk):

take natural log of both sides! for the second one, simplify it a bit before taking log.

OpenStudy (mathmagician):

x-y=2

OpenStudy (lgbasallote):

@lgbasallote

OpenStudy (anonymous):

@Yahoo! , @apoorvk is correct, its a bit like that last question i helped you with

OpenStudy (experimentx):

try to find the value of 5^x and 4^y first ... not sure though..

OpenStudy (anonymous):

the answer is 2

OpenStudy (experimentx):

let 5^x = a and 4^y=b

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

though there might be some other better method ..

OpenStudy (anonymous):

No @experimentX plz continue let see wat u get plz

OpenStudy (experimentx):

a + b = 3189 a/125 + 4 = 281

OpenStudy (anonymous):

i am wondering what the snap way to do this is i get \[a+b=3189, \frac{1}{125}a+4b=281\]

OpenStudy (anonymous):

but since you are looking only for \(x-y\) maybe there is an easier way

OpenStudy (anonymous):

i got b=64 a=3125

OpenStudy (anonymous):

but i think this does not help

OpenStudy (phi):

it helps a lot

OpenStudy (anonymous):

ok so \(4^y=64\implies y=3\) right?

OpenStudy (experimentx):

still there must be some algebraic trick ... if somebody gets it .. let me know first

OpenStudy (anonymous):

similarly \(5^x=3125\implies ...\)

OpenStudy (anonymous):

maybe this is the best way. i cannot think of a snap one notice it does not use logs

OpenStudy (anonymous):

6

OpenStudy (phi):

5^5= 3125

OpenStudy (anonymous):

actually \(5^5=3125\) and \(5-3=2\)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

thxxxx @phi @satellite73 @experimentX

OpenStudy (experimentx):

yw